will you show me the answer to
4x-5y=5
6x-5y=25

Answer by amarjeeth123(569)

4x-5y=5........equation 1
6x-5y=25.......equation 2
Subtracting equation 1 from equation 2 we get,
6x-4x=25-5
2x=20
x=10
Substituting in equation 1 we get,
40-5y=5
5y=35
y=7
The solution is x=10 and y=7.